# Deep Learning Framework Testing via Heuristic Guidance Based on Multiple Model Measurements

### v1

@@ 3.2.3 Model Execution Time Measurement
The execution time of a model, denoted as $time$, is a measurement recorded during test runs. The model execution time is the primary time overhead for each round of framework testing, reflecting the efficiency of the framework testing. Generally, the longer the execution time of a single model, the longer each round of testing will take, and the fewer test input models are executed. Some models that may trigger DL framework bugs may lose the opportunity to be executed, thereby reducing the effectiveness and efficiency of the method. Due to the lack of awareness of the correlation between the bug detection effectiveness of a single model and the model execution time, the execution time is overlooked in existing DL framework testing methods.
3.3 Research Questions
@@
3.4 Empirical Results for RQ1 & RQ2
The Pearson correlation coefficient among model measurements is shown in Table I. The first column represents the method names. The second column shows the Pearson correlation coefficient between operator combination variety and bug detection effectiveness. The third column shows the Pearson correlation coefficient between operator combination variety and model execution time. It is worth noting that LEMON generates only four unique models in six hours, which is too small to calculate Pearson correlation coefficient. Therefore, we mark the empirical results of LEMON as NaN.
TABLE I: Pearson Correlation Coefficient among Model Measurements
@@ 3.4 Empirical Results for RQ1 & RQ2
Muffin 0.21 0.50
COMET 0.32 0.42
Gandalf 0.09 0.48
@@
Finding 1 (Answer to RQ1): Higher operator combination variety contributes to increasing bug detection effectiveness of a model. In the second column of Table I, the Pearson correlation coefficient between operator combination variety and bug detection effectiveness exceeds 0.2 in Muffin and COMET, which indicates a positive correlation between the above two model measurements. Noting that the Pearson correlation coefficient in Gandalf (0.09) does not exceed 0.2, we analyze the models generated by Gandalf and find that these models are too small (with an average of only 5.60 operators per model). As a result, the variation in bug detection effectiveness of the models generated by Gandalf is minimal, and thus the correlation in Gandalf cannot be adequately captured. In conclusion, higher operator combination variety contributes to improving the bug detection effectiveness.
Finding 2 (Answer to RQ2): Higher operator combination variety contributes to increasing the model execution time of a model. In the last column of Table I, the Pearson correlation coefficient between model execution time and operator combination variety exceeds 0.2 in all methods, which indicates a positive correlation between model execution time and operator combination variety. This correlation validates our intuition: the higher the operator combination variety, the more complex the computation performed by the model, and the longer the model execution time.
3.5 Empirical Results for RQ3
To investigate the effect of model execution time on DL framework testing, we sort the models generated by each baseline in descending order of model execution time and divide these models into two groups. One group consists of models with execution times in the higher 50% of all models generated by a baseline, named “larger model”. The other group consists of models with execution times in the lower 50%, named “smaller model”. We record the total execution time for each group, the average execution time per model in each group, the total number of bugs detected by each group, and the average time to detect each bug. The results for COMET and Gandalf are shown in Table II and Table III, respectively. Since LEMON only detects one DL framework bug and Muffin does not detect any DL framework bugs, we exclude them in RQ3.
TABLE II: Effect of Model Execution Time in COMET
@@ 3.5 Empirical Results for RQ3
TABLE III: Effect of Model Execution Time in Gandalf
@@
Finding 3 (Answer to RQ3): Longer model execution time limits the efficiency of DL framework testing. In table II and III, although the total bug number triggered by larger models are more than that triggered by smaller models (12 greater than 7 in COMET, 24 greater than 23 in Gandalf), the bug detection efficiency of smaller models exceeds that of larger models. Specifically, in COMET, smaller models take an average of 993.58 seconds to detect a bug, less than the 1195.41 seconds taken by larger models. In Gandalf, smaller models take an average of 110.38 seconds to detect a bug, less than the 794.21 seconds taken by larger models. The advantage of smaller models in bug detection efficiency comes from their shorter execution time (92.73 seconds shorter than 195.27 seconds in COMET, 8.70 seconds shorter than 65.06 seconds in Gandalf). In conclusion, longer model execution time limits the bug detection efficiency.
4 METHODOLOGY
4.1 Overview
@@
Figure 1: Workflow of DLMMM
4.2 Test Input Tensor Generation
Tensor is an important component of test inputs in DL framework testing. Following existing methods (e.g., LEMON [11]), DLMMM designs two test input tensor generation modes. In one mode, DLMMM loads tensors from publicly available datasets. DLMMM supports all datasets adopted by existing methods, which include MNIST, Fashion-MNIST, CIFAR-10, ImageNet, Sine-Wave, and Stock-Price. In another mode, DLMMM randomly generates test input tensors based on the tensor shape. The tensor shape is specified before testing begins, which is in the format of $NCHW$ ($batch$, $channel$, $height$, $width$).
4.3 Seed Model Generation
@@
4.3.1 Model Mutation
Model mutation is used to generate new DL models. Existing methods design mutation rules (e.g., insert operator, remove operator) and heuristically select which mutation rule to apply. These methods fail to heuristically control the specific implementation of each mutation rule. To overcome this limitation, DLMMM designs a double-level heuristic mutation process based on the DAG representation of the model (introduced in Section 2.2). In DLMMM’s mutation process, mutation rules are implemented as replacing edges in the DAG (e.g., replacing an edge representing the operator $None$ with the edge representing another operator corresponds to the mutation rule “insert operator” in existing methods). By the following heuristic mutation process, in DLMMM, the heuristic indicator not only controls the probability of selecting each mutation rule, but also controls the specific implementation of each mutation (e.g., the probability of inserting or removing each type of operator). The detailed mutation process is as follows. ➀ Randomly sample a predecessor vertex $i$ ($i$ is not the sink). ➁ Randomly sample a successor vertex $j$ ($j>i$). ➂ Sample an operator $o$ based on the $weight$ of each operator. The calculation of $weight$ is introduced in Section 4.6. The probability $p$ of each operator being sampled is:
@@ 4.3.1 Model Mutation
4.3.2 Mutation-Based Seed Model Generation
This process is designed to generate seed models based on the pre-specified model configuration, which determines the number of operators. The detailed process is as follows: ➀ Initialize an empty seed model pool. ➁ Generate a trivial model and add it to the seed model pool. The DAG representation for this trivial model is an operator chain from the source vertex to the sink vertex composed of the operator $identity$. The number of operators in the trivial model is determined in the model configuration. ➂ Randomly select a seed model from the seed model pool, mutate it according to the mutation described in Section 4.3.1, and add the newly generated model into the seed model pool. ➃ Repeat Step 3 until the seed model pool reaches a pre-specified size (the size is configured before testing begins).
4.4 Test Input Model Generation
DLMMM generates new test input models by mutating seed models in the seed model pool (constructed in Section 4.3). In DLMMM, seed model selection and model mutation are both heuristically guided by the heuristic indicator $fitness$, which is calculated based on the recorded model measurements (will be introduced in Section 4.6). In seed model selection, DLMMM uses a tournament algorithm [19] to select $k$ seed models from the seed model pool based on their $fitness$. $K$ is the specified before DLMMM begins, which is usually set to 1 to promote algorithm’s efficiency. The tournament algorithm is performed as follows: ➀ Initialize an empty winner seed model pool. ➁ Randomly sample $k$ seed models from the seed model pool. ➂ Select the seed model with the highest $fitness$ from the subset selected in Step ➁. ➃ Add the selected seed model to the winner seed model pool and remove it from the seed model pool to avoid reselection. ➄ Repeat Step ➁ - ➃ until $k$ models are added into the winner seed model pool. The winner seed model pool is the output of the tournament algorithm.
@@ 4.4 Test Input Model Generation
4.5 Differential Testing
DLMMM adopts differential testing among different DL frameworks. The test inputs include the test input tensors generated in Section 4.2 and the test input models generated in Section 4.4. DLMMM executes the above test inputs under different DL frameworks and records execution logs and output tensors. By analyzing logs and output tensors, DLMMM targets two types of bugs: non-numerical bugs and numerical bugs. Specifically, the non-numerical bugs include crashes. The numerical bugs include Not-A-Number (NaN) bugs, and inconsistency bugs.
@@ 4.5 Differential Testing
NaN & inconsistency bugs. DLMMM designs a voting mechanism to figure out which framework triggers NaN & inconsistency bugs. Specifically, NaN bugs are detected when the output tensor of a framework contains NaN while the output tensors of other frameworks do not. Inconsistency bugs are detected when two $inconsistency$ (see Section 3.2.1) calculated based on the output tensor under the same framework both exceed the threshold $\epsilon$. When more than two $inconsistency$ exceed $\epsilon$, DLMMM attributes this inconsistency to the common framework related to the largest two $inconsistency$. To ensure fairness in the experiment and better eliminate errors caused by computational randomness, DLMMM sets the value of the threshold to the maximum value among all existing methods (e.g., Predoo [20], Gandalf [15], Muffin [12], etc). The value of the threshold $\epsilon$ is 0.15. In addition, a NaN & inconsistency bug may be triggered by multiple same models, which leads to redundancy. To avoid redundancy, DLMMM analyzes the model structures and removes redundant models.
@@
4.6 Heuristic Guidance
Heuristic guidance is an important component in DL framework testing, which is usually used to control test input model generation. As found in Section 3, there are correlations between multiple model measurements. To achieve a trade-off between multiple model measurements and improve the overall testing effectiveness, DLMMM designs a measurement fusion based on the correlations between model measurements. Specifically, DLMMM fuse bug detection effectiveness, operator combination variety, and model execution time into the heuristic indicator $fitness$, which is expected to meet the following five requirements. ➀ $Fitness$ should reflect the overall level of the models in bug detection effectiveness, model execution time and operator combination variety. ➁ $Fitness$ should exhibit positive correlation with bug detection effectiveness and operator combination variety, and should exhibit negative correlation with model execution time. ➂ $Fitness$ should capture the contrast intensity in each measurement among different models. ➃ $Fitness$ should capture the correlation between different measurements ➄ $Fitness$ should eliminate dimensional differences among different measurements.
To meet the above five requirements, DLMMM’s measurement fusion is designed based on CRITIC [21]. CRITIC is a classical objective weighting method, which is used to determine the objective weights of all measurements based on their contrast intensity and conflict. Specifically, contrast intensity is calculated by the standard deviation, and conflict between measurements is based on their correlation. In DLMMM, the $fitness$ of a newly generated model is calculated based on the model measurements of itself and all models generated before, which is recorded in differential testing (see Section 4.5). For convenience in calculation, we place all recorded model measurements into a matrix named $judge\_matrix$. The rows in the $judge\_matrix$ represent the measurement values of each model. The first column in the $judge\_matrix$ represents bug detection effectiveness, denoted as $performance\_column$. The second column represents operator combination variety, denoted as $variety\_column$. The third column represents model execution time, denoted as $time\_column$. The detailed calculation for $fitness$ is as follows:
➀ Replace each value in $time\_column$ by its reciprocal to maintain a negative correlation between model execution time and the fused heuristic indicator.
➁ Normalize $performance\_column$, $variety\_column$, and $time\_column$ as follows:
@@ 4.6 Heuristic Guidance
In this formula, the meanings of $q$ are the same as in Step 2. $r_{q}$ is the Pearson correlation coefficient, which reflects the correlation between measurements.
@@
➆ Calculate the heuristic indicator $fitness$ of the newly generated model based on its original data $x_{q}$ in the $judge\_matrix$ and the weights of each column $w_{q}$ (calculated in Step 6) as follows:
@@
The $fitness$ calculated using the above measurement fusion depicts the contrast intensity of the data from different models within each model measurement, and the correlation between multiple measurements. In addition, the above measurement fusion eliminates the dimensional differences between different model measurements. Therefore, $fitness$ reflects the overall level of generated models in bug detection effectiveness, operator combination variety, and model execution time. Based on the $fitness$ of the seed model and the newly generated model, DLMMM calculates the contribution of this mutation $\Delta fitness$ as follows:
@@
where $fitness_{new}$ is the $fitness$ of the newly generated model and $fitness_{seed}$ is the $fitness$ of the seed model. Based on $\Delta fitness$, DLMMM updates the $weight$ of each operator, which is used to calculate the probability being sampled in mutation. The update formula is as follows:
@@
where $weight_{1}$ and $weight_{0}$ is the $weight$ after and before updating, respectively.
5 Evaluation
To evaluate the effectiveness of DLMMM in detecting DL framework bugs, we compare DLMMM with four baselines: LEMON [11], Muffin [12], COMET [13], and Gandalf [15]. Because there are no publicly available codes for graphFuzz [17], we do not select it as our baseline. All codes and experimental results are available in our public repository.
5.1 Experimental Setup
We test three popular DL frameworks: TensorFlow [22], PyTorch [23], and MindSpore [24]. The default software versions are: TensorFlow v2.9.0, PyTorch v1.12.0, and MindSpore v2.1.0. The workstation is equipped with the Ubuntu 22.04 operating system and an Intel Core Processor (Skylake) (16cores, 2.0GHz). The experimental time on each dataset (see Section 4.2) is six hours. The threshold $\epsilon$ for detecting inconsistency bugs is 0.15 (see Section 4.5).
5.2 Research Questions

### v2

@@ 3.2.3. Model Execution Time Measurement
The execution time of a model, denoted as timetime, is a measurement recorded during test runs. The model execution time is the primary time overhead for each round of framework testing, reflecting the efficiency of the framework testing. Generally, the longer the execution time of a single model, the longer the time each round of testing takes, and the fewer test input models are executed. Some models that may trigger DL framework bugs may lose the opportunity to be executed, thereby reducing the effectiveness and efficiency of the method. Due to the lack of awareness of the correlation between the bug detection effectiveness of a single model and the model execution time, the execution time is overlooked in existing DL framework testing methods.
3.3. Research Questions
@@
3.4. Empirical Results for RQ1 & RQ2
The Spearman correlation coefficient among model measurements is shown in Table 1. The first column represents the method names. The second column shows the Spearman correlation coefficient between the operator combination variety and bug detection effectiveness. The third column shows the Spearman correlation coefficient between the operator combination variety and model execution time. It is worth noting that LEMON generates only four unique models in six hours, which is too small to calculate Spearman correlation coefficient. Therefore, we mark the empirical results of LEMON as NaN.
Table 1. Spearman Correlation Coefficient among Model Measurements
@@ 3.4. Empirical Results for RQ1 & RQ2
Muffin 0.43 0.56
COMET 0.44 0.38
Gandalf 0.52 0.79
@@
Finding 1 (Answer to RQ1): Higher operator combination variety contributes to increasing bug detection effectiveness of a model. In the second column of Table 1, the Spearman correlation coefficient between operator combination variety and bug detection effectiveness exceeds 0.2 in all methods (0.43 in Muffin, 0.44 in COMET, and 0.52 in Gandalf), which indicates a positive correlation between the above two model measurements. In conclusion, higher operator combination variety contributes to improving the bug detection effectiveness.
Finding 2 (Answer to RQ2): Higher operator combination variety contributes to increasing the model execution time of a model. In the last column of Table 1, the Spearman correlation coefficient between model execution time and operator combination variety exceeds 0.2 in all methods (0.56 in Muffin, 0.38 in COMET, and 0.79 in Gandalf), which indicates a positive correlation between model execution time and operator combination variety. This correlation validates our intuition: the higher the operator combination variety, the more complex the computation performed by the model, and the longer the model execution time.
3.5. Empirical Results for RQ3
To investigate the effect of model execution time on DL framework testing, we sort the models generated by each baseline in descending order of model execution time and divide these models into two groups. One group consists of models with execution times in the higher 50% of all models generated by a baseline, named “larger model”. The other group consists of models with execution times in the lower 50%, named “smaller model”. We record the total execution time for each group, the average execution time per model in each group, the total number of bugs detected by each group, and the average time to detect each bug. The results for COMET and Gandalf are shown in Table 2 and Table 3, respectively. Since LEMON only detects one DL framework bug and Muffin does not detect any DL framework bugs, we exclude them in RQ3.
Table 2. Effect of Model Execution Time in COMET
@@ 3.5. Empirical Results for RQ3
Table 3. Effect of Model Execution Time in Gandalf
@@
Finding 3 (Answer to RQ3): Longer model execution time limits the efficiency of DL framework testing. In table 2 and 3, although the total bug number triggered by larger models are more than that triggered by smaller models (12 greater than 7 in COMET, 24 greater than 23 in Gandalf), the bug detection efficiency of smaller models exceeds that of larger models. Specifically, in COMET, smaller models take an average of 993.58 seconds to detect a bug, less than the 1195.41 seconds taken by larger models. In Gandalf, smaller models take an average of 110.38 seconds to detect a bug, less than the 794.21 seconds taken by larger models. The advantage of smaller models in bug detection efficiency comes from their shorter execution time (92.73 seconds shorter than 195.27 seconds in COMET, 8.70 seconds shorter than 65.06 seconds in Gandalf). In conclusion, longer model execution time limits the bug detection efficiency.
4. METHODOLOGY
4.1. Overview
@@
Figure 1. Workflow of DLMMM
4.2. Test Input Tensor Generation
Tensor is an important component of test inputs in DL framework testing. Following existing methods (e.g., LEMON (lemon, )), DLMMM designs two test input tensor generation modes. In one mode, DLMMM loads tensors from publicly available datasets. DLMMM supports all datasets adopted by existing methods, which include MNIST, Fashion-MNIST, CIFAR-10, ImageNet, Sine-Wave, and Stock-Price. In another mode, DLMMM randomly generates test input tensors based on the tensor shape. The tensor shape is specified before testing begins, which is in the format of NCHWNCHW (batchbatch, channelchannel, heightheight, widthwidth).
4.3. Seed Model Generation
@@
4.3.1. Model Mutation
Model mutation is used to generate new DL models. Existing methods design mutation rules (e.g., insert operator, remove operator) and heuristically select which mutation rule to apply. These methods fail to heuristically control the specific implementation of each mutation rule. To overcome this limitation, DLMMM designs a double-level heuristic mutation process based on the DAG representation of the model (introduced in Section 2.2). In DLMMM’s mutation process, mutation rules are implemented as replacing edges in the DAG (e.g., replacing an edge representing the operator $None$ with the edge representing another operator corresponds to the mutation rule “insert operator” in existing methods). By the following heuristic mutation process, in DLMMM, the heuristic indicator not only controls the probability of selecting each mutation rule, but also controls the specific implementation of each mutation (e.g., the probability of inserting or removing each type of operator). The detailed mutation process is as follows. ➀ Randomly sample a predecessor vertex $i$ (ii is not the sink). ➁ Randomly sample a successor vertex $j$ ($j>i$). ➂ Sample an operator $o$ based on the weightweight of each operator. The calculation of weightweight is introduced in Section 4.6. The probability $p$ of each operator being sampled is:
@@ 4.3.1. Model Mutation
4.3.2. Mutation-Based Seed Model Generation
DLMMM generates test input models from sketch. Specifically, before the method begins, the pre-specified model configuration determines the number of operators contained in each model. After that, DLMMM firstly initializes an empty seed model pool, and generates an seed model without any operators, which is called the trivial model. The DAG representation for this trivial model is an operator chain from the source vertex to the sink vertex composed of the operator identityidentity. Then, DLMMM randomly selects a seed model from the seed model pool, mutate it according to the mutation described in Section 4.3.1, and add the newly generated model into the seed model pool. Finally, DLMMM repeats the above model selection and mutation until the seed model pool reaches a pre-specified size (configured before testing begins).
4.4. Test Input Model Generation
DLMMM generates new test input models by mutating seed models in the seed model pool (constructed in Section 4.3). In DLMMM, seed model selection and model mutation are both heuristically guided by the heuristic indicator fitnessfitness, which is calculated based on the recorded model measurements (will be introduced in Section 4.6). In seed model selection, DLMMM uses a tournament algorithm (tournament, ) to select $k$ seed models from the seed model pool based on their fitnessfitness. $K$ is the specified before DLMMM begins, which is usually set to 1 to promote algorithm’s efficiency. The tournament algorithm is performed as follows. ➀ Initialize an empty winner seed model pool. ➁ Randomly sample $k$ seed models from the seed model pool. ➂ Select the seed model with the highest fitnessfitness from the subset selected in Step ➁. ➃ Add the selected seed model to the winner seed model pool and remove it from the seed model pool to avoid reselection. ➄ Repeat Step ➁ - ➃ until $k$ models are added into the winner seed model pool. The winner seed model pool is the output of the tournament algorithm.
@@ 4.4. Test Input Model Generation
4.5. Differential Testing
DLMMM adopts differential testing among different DL frameworks. The test inputs include the test input tensors generated in Section 4.2 and the test input models generated in Section 4.4. DLMMM executes the above test inputs under different DL frameworks and records execution logs and output tensors. By analyzing logs and output tensors, DLMMM targets both non-numerical bugs and numerical bugs. Specifically, the non-numerical bugs include crashes. The numerical bugs include Not-A-Number (NaN) bugs, and inconsistency bugs.
@@ 4.5. Differential Testing
NaN & inconsistency bugs. DLMMM designs a voting mechanism to figure out which framework triggers NaN & inconsistency bugs. Specifically, NaN bugs are detected when the output tensor of a framework contains NaN while the output tensors of other frameworks do not. Inconsistency bugs are detected when two inconsistencyinconsistency (see Section 3.2.1) calculated based on the output tensor under the same framework both exceed the threshold $\epsilon$. When more than two inconsistencyinconsistency exceed $\epsilon$, DLMMM attributes this inconsistency bug to the common framework related to the largest two inconsistencyinconsistency. To ensure fairness in the experiment and better eliminate errors caused by computational randomness, DLMMM sets the value of the threshold to the maximum value among all existing methods (e.g., Predoo (predoo, ), Gandalf (gandalf, ), and Muffin (muffin, )). The value of the threshold $\epsilon$ is 0.15. In addition, an NaN & inconsistency bug may be triggered by multiple the same models, which leads to redundancy. To avoid redundancy, DLMMM analyzes the model structures and removes redundant models.
@@
4.6. Heuristic Guidance
Heuristic guidance is an important component in DL framework testing, which is usually used to control test input model generation. As found in Section 3, there are correlations between multiple model measurements. To achieve a trade-off between multiple model measurements and improve the overall testing effectiveness, DLMMM designs a measurement fusion based on the correlations between model measurements. Specifically, DLMMM fuse bug detection effectiveness, operator combination variety, and model execution time into the heuristic indicator fitnessfitness, which is expected to meet the following five requirements. ➀ $Fitness$ should reflect the overall level of the models in bug detection effectiveness, model execution time and operator combination variety. ➁ $Fitness$ should exhibit positive correlation with bug detection effectiveness and operator combination variety, and should exhibit negative correlation with model execution time. ➂ $Fitness$ should capture the contrast intensity in each measurement among different models. ➃ $Fitness$ should capture the correlation between different measurements ➄ $Fitness$ should eliminate dimensional differences among different measurements.
To meet the above five requirements, DLMMM’s measurement fusion is designed based on CRITIC (CRITIC, ). CRITIC is a classical objective weighting method, which is used to determine the objective weights of all measurements based on their contrast intensity and conflict. Specifically, contrast intensity is calculated by the standard deviation, and conflict between measurements is based on their correlation. In DLMMM, the fitnessfitness of a newly generated model is calculated based on the model measurements of itself and all models generated before, which is recorded in differential testing (see Section 4.5). For convenience in calculation, we place all recorded model measurements into a matrix named judge_matrixjudge\_matrix. The rows in the judge_matrixjudge\_matrix represent the measurement values of each model. The first column in the judge_matrixjudge\_matrix represents bug detection effectiveness, denoted as performance_columnperformance\_column. The second column represents operator combination variety, denoted as variety_columnvariety\_column. The third column represents model execution time, denoted as time_columntime\_column. The detailed calculation for fitnessfitness is as follows:
➀Replace each value in time_columntime\_column by its reciprocal to maintain a negative correlation between model execution time and the fused heuristic indicator.
➁Normalize performance_columnperformance\_column, variety_columnvariety\_column, and time_columntime\_column as follows:
@@ 4.6. Heuristic Guidance
In this formula, the meanings of $q$ are the same as in Step 2. $r_{q}$ is the Spearman correlation coefficient, which reflects the correlation between measurements.
@@
➆Calculate the heuristic indicator fitnessfitness of the newly generated model based on its original data $x_{q}$ in the judge_matrixjudge\_matrix and the weights of each column $w_{q}$ (calculated in Step 6) as follows:
@@
The fitnessfitness calculated using the above measurement fusion depicts the contrast intensity of the data from different models within each model measurement, and the correlation between multiple measurements. In addition, the above measurement fusion eliminates the dimensional differences between different model measurements. Therefore, fitnessfitness reflects the overall performance of generated models in bug detection effectiveness, operator combination variety, and model execution time. Based on the fitnessfitness of the seed model and the newly generated model, DLMMM calculates the contribution of this mutation $\Delta fitness$ as follows:
@@
where $fitness_{new}$ is the fitnessfitness of the newly generated model and $fitness_{seed}$ is the fitnessfitness of the seed model. Based on $\Delta fitness$, DLMMM updates the weightweight of each operator, which is used to calculate the probability being sampled in mutation. The update formula is as follows:
@@
where $weight_{1}$ and $weight_{0}$ is the weightweight after and before updating, respectively.
5. Evaluation
To evaluate the effectiveness of DLMMM in detecting DL framework bugs, we compare DLMMM with four baselines: LEMON (lemon, ), Muffin (muffin, ), COMET (comet, ), and Gandalf (gandalf, ). Because there are no publicly available codes for graphFuzz (fanggraphbased, ), we do not select it as our baseline. All codes and experimental results are available in our public repository.
5.1. Experimental Setup
We test three popular DL frameworks: TensorFlow (tensorflow, ), PyTorch (pytorch, ), and MindSpore (mindspore, ). The default software versions are: TensorFlow v2.9.0, PyTorch v1.12.0, and MindSpore v2.1.0. The workstation is equipped with the Ubuntu 22.04 operating system and NVIDIA RTX 4090D GPU, and an Intel Core Processor (Skylake) (16cores, 2.0GHz). The experimental time on each dataset (see Section 4.2) is six hours. No new bugs are detected during the last hour of each experiment, demonstrating the convergence. The threshold $\epsilon$ for detecting inconsistency bugs is 0.15 (see Section 4.5).
5.2. Research Questions